OpenStudy (anonymous):

Solve the Equation. (√2x+15)=(√x+10)

OpenStudy (anonymous):

(√2x=15)=(√x+10)??? typo?

OpenStudy (anonymous):

Made a correction

hartnn (hartnn):

square both sides, what do u get?

OpenStudy (anonymous):

2x+15=x+10

hartnn (hartnn):

good! now subtract x from both sides: )

OpenStudy (anonymous):

x+15=10 x=-5